What is 10 to the 2/3 times 1,000 equal?
Answer by htmentor(1343)

Since 1000 = 10^3, we can write the expression as
1000*10^(2/3) = 10^3*10^(2/3)
Since these two terms have the same base, we can add the exponents together
10^3*10^(2/3) = 10^(3+2/3) = 10^(11/3)